what is the answer to -7(z+3)+4(3z-4)=4(z-3)+9?
Answer by checkley79(3341)

-7(z+3)+4(3z-4)=4(z-3)+9
-7Z-21+12Z-16=4Z-12+9
5Z-4Z=-3+37
Z=34 ANS.
PROOF:
-7(34+3)+4(3*34-4)=4(34-3)+9
-7*37+4(102-4)=4*31+9
-259+4(98)=124+9
-259+392=133
133=133